NASA MIR Program Mir 24 / NASA 6 Status

NASA MIR Program Mir 24 / NASA 6 Status Date:09
Wednesday, December 17, 1997 Mission Day:09
Mir 24 / NASA 6 MD 134/83

Mir Station System/Subsystem Activities All primary major systems continue to operate nominally. Inspector activities were terminated today due to a problem with the Inspector star tracker. The crew kept receiving an input that the star tracker was overheating. The logic of the Inspector payload is such that without an operating start tracker, no burns can be performed

All Progress burns were executed with no anomalies. The Inspector and the Progress are on safe trajectories and there are no trajectory concerns. The crew did not perform burns to separate Mir from the Inspektor. According to Shift Flight Director Viktor Shedrin, there was a formatting problem in software uplinked to the Mir, and this prevented the burn from occurring. The current plan is to wait until the morning, when the crew will determine the distance to Inspektor, and a decision will be made at that time whether to reattempt a station burn. At about 21:00 DMT, the crew stated that the Inspektor was approximately 425 meters behind and 45-47 degrees higher than the Mir. The entire crew went to sleep following the comm pass ending at 21:20.

Payload

ACTIVITIES FB:HSR:I:230909CHAPAT Experiment Computer Monitor MG:BT:I:8 10909IPCG Growth Rate Status Check HLS:BP:I:100909

Mir Crew: Completion of Computerized Questionnaire Packet HLS:BP:I:130909NASA Crew: C ompletion of Computerized Questionnaire Packet HLS:MS:I:55R09Calcium K inetics Protocol - Blood/Saliva Collection HLS:MS:I:740909Calcium Kin etics Protocol - 24-hour Urine Collection HLS:MS:I:750909Calcium Kine tics Protocol - FFME Log Prepack/EVA prep OTHER ACTIVITIES Inspe ctor operations.

Data Management On-Board -nothing to report.

Network Status -nothing to report. Ground Facilities Status - not hing to report. Operations Procedures/Timelines The timelines for W ednesday, December 17, and Friday, December 19, were submitted to POSA for distribution. Experiment Operations Bone - Dave did not repor t on continuance of Bone, however we anticipate that activities occurre d as planned because he was aware of what was scheduled. IPCG- Dave st ated that he had begun experiment number one in cell 5. He saw little crystal growth after nucleation, but after turning on power for #1 he o bserved an almost immediate 93boulder-field94 of crystals. While he tr ied to focus on one crystal, he had to focus on part of this 93boulder -field.94 He stated that there was therefore obvious precrystallizati on CHAPAT - Dave tried to report CHAPAT data but static prevented MOST from recording this data. Experiment Hardware Nothing to report

Radiograms A Radiogram for BSTC Locker Reconfig has been submitt ed Stowage/Transfer Dave reported that he has received the pre-pac k list.

SMP/Medical A PMC occurred from 19:58 - 20:02 DMT.

En vironmental Parameters (last 24 hours): Parameter09Min.09Max.09Nomin al Range Pressure (mmHg) Base Block096980970609660 - 850 ppO2 (mm Hg)091600916209140 - 200 ppCO2 (mmHg)094.3094.609up to 8.0 ppH2 0 (mmHg)0913.70914.309 8 - 14 Humidity (percentage)094709530930 - 75 Temperature (B0C) base block (BB)0927.40928.80918 - 28

Co mmunications Summary A/G with Consultant Group Due to Inspector O perations, no A/G with Consultant Group was scheduled today. Packet One work file (English form 24) was submitted for uplink.

  • submit to reddit


Loading



Humans to Mars Summit 2019
The Robert H. Goddard Memorial Symposium
Military Space Situational Awareness April 1 - April 2, 2019, London, UK
Aster Labs - CubeSats




Join our mailing list




Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on December 17, 1997 12:17 PM.

Read Dan Goldin's lips: Better-Cheaper-Faster is Better-Cheaper-Faster was the previous entry in this blog.

Mir Status Report 23 December 1997 is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.